# Synergizing LLMs and Knowledge Graphs: A Novel Approach to Software Repository-Related Question Answering

### v1

@@ 5. Results
In this section, we first present the results of the exploratory analysis, and then present the results of the three main research questions. For each research question, we present the motivation, the approach, and the results.
@@ 5.1. RQ0: How good are LLMs in generating Cypher queries for knowledge graphs?
$\mathbbm{1}(V_{n},\hat{V}_{n})=\begin{cases}1,&\text{if }V_{n}=\hat{V}_{n}\\ 0,&\text{if }V_{n}\neq\hat{V}_{n}\end{cases}$
@@ 5.2. RQ1: How effective is our approach in answering software repository-related questions?
For this purpose, we evaluated the 20 questions described in Section 4.2 for each of the selected projects. For each question, we also executed the process five times to account for the stochastic nature of LLMs in the generation process. We compared the final responses generated by our approach to the oracle answers (pre-determined correct answers based on the data in the knowledge graph). A question was considered correctly answered if our approach provided the correct response at least 50% of the time. If it fails in three or more executions, the question is marked as incorrect.
@@
The limitation on incorrectly modeling relationships, faulty arithmetic logic, misapplied attribute filtering, and date formatting highlights the limitation in the reasoning ability of the LLM. This hinders the LLM in accurately interpreting and utilizing the right nodes and relationships within the knowledge graph. This aligns with prior studies that highlight the challenges LLMs face in reasoning tasks (Chen et al., 2024; Wang et al., 2023a, b). The incorrect reasoning often leads to misconstructed queries that do not align with the intended question, resulting in inaccurate or irrelevant answers. Addressing these limitations is crucial for improving our approach. By refining the LLM’s understanding of the knowledge graph schema and enhancing its reasoning capabilities, we can reduce the incidence of incorrect responses.
@@
Incorrect relationship modeling The LLM deviates from the intended logic of the question, leading to a query that incorrectly models relationships within the knowledge graph 123 75.0%
@@
Motivation. In RQ2, we identified that the faulty reasoning of the LLM during query generation is the main contributor negatively affecting the accuracy of our approach in answering the questions. Prior studies have shown that using chain-of-thought to generate a series of intermediate steps before the final answer can improve the reasoning ability of LLMs (Kojima et al., 2022; Wei et al., 2022). This can be achieved under zero-shot settings (prompting the LLM to think step by step) (Kojima et al., 2022) and few-shot settings (prompting the LLMs with a few chain-of-thought examples). In this research question, our goal is to improve the reasoning ability of our approach to mitigate faulty reasoning. We do this by evaluating our approach using the chain-of-thought prompting by feeding the LLM with step-by-step reasoning examples to guide it in generating a reasoning path to reach an answer.
@@
Approach. To evaluate if chain-of-thought prompting can improve the performance of our approach in answering software repository questions, we first evaluated the approach by including a zero-shot chain-of-thought instruction, that is: “Let’s think step by step” (Kojima et al., 2022) to the prompt template in Figure 3. This did not improve the results presented in RQ1 (See Appendix D for the results of the zero-shot chain-of-thought across the selected projects and difficulty levels). We experimented with few-shot chain-of-thought by incorporating examples into the prompt provided to the Query Generator. Similar to (Wei et al., 2022), we adopted the format that begins with the input question, the step-by-step reasoning process, and the final output in our few-shot chain-of-thought prompting. This format aims to guide the LLM in generating intermediate reasoning steps before arriving at the final answer. We constructed the chain-of-thought prompt as presented in Figure 5. The prompt consists of two examples consisting of difficulty level 2 and 3 questions. By providing these examples, we intended to show the LLM how to generate reasoning paths that can help in constructing correct queries. It is important to note that the questions used in the chain-of-thought examples are not part of our evaluation questions.
@@
For the evaluation, we used the same set of 20 questions described in Section 4 and used in RQ1. Also, we followed a similar approach as the evaluation of RQ1; executing the experiments five times for each question to account for the stochastic nature of the LLMs generation (Edge et al., 2024). A question is considered correctly answered if the correct response is generated most of the time otherwise, it is marked as incorrect.
@@
The chain-of-thought prompting approach answered the repository-related question 84% of the time, up from an initial 65% without chain-of-thought. Specifically, the accuracy of the complex questions requiring two or more relationships increased from 50% to 90%. This implies that chain-of-thought can help answer complex questions requiring multiple relationships.
@@ 6. Discussions
In this study, we explored the synergy between large language models (LLMs) and knowledge graphs to enhance the accuracy of software engineering chatbots in answering software repository-related questions. Our findings demonstrate that synergizing LLMs and knowledge graphs can effectively improve the performance of LLM-based chatbots in answering software repositories-related questions.
We evaluated our approach on five popular open-source GitHub projects, with an accuracy range of 60%–75% in RQ1 and 80%–90% in RQ3. The difference in the accuracy of the approach on the projects is due to the non-deterministic nature of the LLM rather than the information in the projects. Our approach relies on the schema of the knowledge graph (see Figure 2), which is the same for all evaluated projects, to generate a query for retrieving the relevant data. Thus, if the LLM generated the same Cypher query for each question each time the question was executed, the accuracy would be constant across the repositories. This shows that the approach is generalizable across projects.
We identified faulty reasoning as the prevalent challenge facing the approach, accounting for four of the limitations. This highlights the LLM’s challenges in accurately interpreting complex relationships and reasoning over the knowledge graph. In addressing the reasoning limitations, we introduced few-shot chain-of-thought prompting. This technique significantly improved the chatbot’s performance by 19% percentage points compared to the initial accuracy.
@@
Our approach achieved results beyond the evaluation questions in Section 4. Our approach was able to answer ad-hoc queries that were not part of the original test set. For example, using lines of code as a metric for productivity (Maxwell et al., 1996; Cheng et al., 2022), we asked the question “Who added the most lines of code in December 2023” on the Vue project, it responded correctly by going through the right reasoning (See Appendix F). This highlights the promising potential of using LLMs with knowledge graphs to transform software engineering chatbots, making them more capable of handling a wide range of user queries.
@@ 7.3. External Validity
External validity concerns the generalizability of our findings beyond the context of this study. A threat to external validity is the generalization beyond the evaluation questions used in our study. We evaluated our approach using a set of 20 questions representing 10 intents described in Section 4. Although these questions cover different difficulty levels and intents, they may not encompass the full diversity of questions that users might pose in real-world scenarios. The chatbot’s performance on these questions might not generalize to other questions, especially those that are more complex and ambiguous. To address this threat, we posed ad-hoc questions to our approach and observed that the approach was able to answer the ad-hoc questions beyond the evaluation set. This suggests some level of generalizability. However, there is still a possibility that applying our approach to different types of questions could yield different results.
@@ 8. Conclusion
In this study, we investigated the synergy between large language models (LLMs) and knowledge graphs to improve the accuracy of software engineering chatbots in answering software repository-related questions. Our approach aimed to accurately answer natural language questions by generating Cypher queries to retrieve relevant repository data from the knowledge graph. Then use the retrieved information as context to generate a natural language response, making repository information accessible to both technical and non-technical stakeholders. We empirically evaluated our approach using five popular open-source GitHub repositories and a set of 20 questions curated from Abdellatif et al. (2020a) and categorized into three levels of difficulties. The findings demonstrated that LLMs, specifically the GPT-4o model, can answer repository-related questions by generating Cypher queries to retrieve accurate data from the knowledge graph. The initial accuracy of 65% achieved by our approach highlighted the potential limitation of synergizing LLMs with knowledge graphs. We manually investigated the instances where the approach failed to generate an accurate response and identified the faulty reasoning by the LLM as the predominant factor (80.5%) affecting the approach. We conducted further empirical evaluation if using few-shot chain-of-thought prompting can improve the accuracy. This technique significantly enhanced the reasoning ability of the LLM in our approach and improved the overall accuracy from 65% to 84%. There was a notable increase in the accuracy of the level 3 questions from 50% to 90%, signifying an improvement in the approach to handling complex queries. Our findings highlight the integration of LLMs with knowledge graphs as a viable solution for making repository data accessible to both technical and non-technical stakeholders. Also, our study highlights the importance of enhancing reasoning capabilities in LLMs. This opens avenues for further investigation in this direction

### v2

@@ 5. Results
In this section, we first present the results of the exploratory analysis, and then present the results of the four main research questions. For each research question, we present the motivation, the approach, and the results.
@@ 5.1. RQ0: How good are LLMs in generating Cypher queries for knowledge graphs?
[…]
@@ 5.2. RQ1: How effective is our approach in answering software repository-related questions?
For this purpose, we evaluated the 20 questions described in Section 4.2 for each of the selected projects. For each question, we also executed the process five times to account for the stochastic nature of LLMs in the generation process. We compared the final responses generated by our approach to the oracle answers (predetermined correct answers based on the data in the knowledge graph). A question was considered correctly answered if our approach provided the correct response in at least 3 out of 5 executions. If it failed in three or more executions, the question was marked as incorrect.
@@
The limitation on incorrectly modeling relationships, faulty arithmetic logic, misapplied attribute filtering, and date formatting highlights the limitation in the reasoning ability of the LLM. This hinders the LLM in accurately interpreting and utilizing the right nodes and relationships within the knowledge graph. This aligns with prior studies that highlight the challenges LLMs face in reasoning tasks (Chen et al., 2024; Wang et al., 2023b, c). The incorrect reasoning often leads to misconstructed queries that do not align with the intended question, resulting in inaccurate or irrelevant answers. Addressing these limitations is crucial for improving our approach. By refining the LLM’s understanding of the knowledge graph schema and enhancing its reasoning capabilities, we can reduce the incidence of incorrect responses.
@@
Incorrect relationship modelling The LLM deviates from the intended logic of the question, leading to a query that incorrectly models relationships within the knowledge graph 123 75.0%
@@
Motivation. In RQ2, the failure analysis showed that most of the limitations leading to an incorrect response are due to the faulty reasoning of the LLM during query generation. To address this limitation, we introduce chain-of-thought (CoT) prompting in the LLM underlying the Query Generator. Our goal for this research question is to improve the reasoning ability of the LLM in the Query Generator to minimize faulty reasoning. We do this by introducing chain-of-thought (CoT) prompting into our approach, feeding the LLM step-by-step reasoning examples to guide it in generating a reasoning path to reach an answer. Prior studies have shown that using chain-of-thought (CoT) to generate a series of intermediate steps before the final answer can improve the reasoning ability of LLMs (Kojima et al., 2022; Wei et al., 2022). This can be achieved under zero-shot settings (prompting the LLM to think step by step) (Kojima et al., 2022) and few-shot settings (prompting the LLM with a few chain-of-thought examples).
@@
Approach. To evaluate if chain-of-thought prompting can improve the performance of our approach in answering software repository questions, we first evaluated the approach by including a zero-shot chain-of-thought instruction, that is: “Let’s think step by step” (Kojima et al., 2022) to the prompt template in Figure 3. This did not improve the results presented in RQ1 (See Appendix D for the results of the zero-shot chain-of-thought across the selected projects and difficulty levels). We experimented with few-shot chain-of-thought by incorporating examples into the prompt provided to the Query Generator. Similar to Wei et al. (2022), we adopted the format that begins with the input question, the step-by-step reasoning process, and the final output in our few-shot chain-of-thought prompting. This format aims to guide the LLM in generating intermediate reasoning steps before arriving at the final answer. We constructed the chain-of-thought prompt as presented in Figure 5. The prompt consists of two examples, consisting of difficulty level 2 and 3 questions. By providing these examples, we intended to show the LLM how to generate reasoning paths that can help in constructing correct queries. It is important to note that the questions used in the chain-of-thought examples are not part of our evaluation questions.
@@
For the evaluation, we first measure the gains in execution accuracy made by incorporating chain-of-thought and then benchmark against the baselines described in Section 4.4. To evaluate the gains, we use the same set of 20 questions described in Section 4 and used in RQ1, executing the experiments five times for each question to account for the stochastic nature of the LLM’s generation (Edge et al., 2024). A question is considered correctly answered if the correct response is generated most of the time; otherwise, it is marked as incorrect.
For the baseline comparison, we use the 150 evaluation questions described in Section 4.2 and run each baseline five times, as done previously. The baselines include the intent-based MSRBot and the GPT-4o model with web search capabilities. MSRBot produces identical answers across runs due to its deterministic intent-matching process, while GPT-4o-search-preview is run with the default parameters through the OpenAI API across iterations to ensure a fair comparison.
@@
Comparison with baselines. We further compared our few-shot chain-of-thought approach with the intent-based MSRBot and the GPT-4o model with web search, which serve as baselines for this task (see Section 4.4). Table 15 summarises the accuracy of the three methods across the five projects. Our approach achieved an average accuracy of 0.82 across 750 evaluation questions, outperforming MSRBot (0.70) and GPT-4o-search-preview (0.19). The GPT-4o-search baseline struggled to answer the repository-related questions, corroborating prior observations that retrieval-augmented LLMs often fail, although this approach outperformed the prior study (Abedu et al., 2024a). Table 16 reports the execution accuracy by question difficulty. Our approach consistently outperformed MSRBot on level-1 and level-3 questions, achieving 0.87 accuracy on the most complex level-3 questions compared to 0.60 for MSRBot. For level-2 questions, MSRBot slightly outperformed our approach. GPT-4o-search performed poorly across all levels, reinforcing that naïve web-enabled LLMs are not yet suitable for this task. Overall, few-shot chain-of-thought prompting yields higher accuracy than the baselines on most question types.
Consistency of responses. Beyond execution accuracy, we examined the consistency of responses across multiple runs of the same question for our approach. Specifically, we counted the number of times the correct answer was generated in the five runs for each question. The distribution in Figure 6 shows this for all five projects. Overall, 76.4% of questions were answered correctly each time the question was repeated (5/5 correct runs). This increases to 81.7% under our majority-vote criterion (i.e., 3/5, 4/5, or 5/5 correct runs) in our evaluation. This shows that our approach is not only accurate on average but also reliably produces the correct answer most of the time. This stability is an important quality for practical deployment, as it reduces variability in responses for the user.
@@
Table 15. Comparison with baselines across the selected projects
Project Questions Correct Accuracy
Our Approach MSRBot GPT-4o- Search-Preview Our Approach MSRBot GPT-4o- Search-Preview
AutoGPT 150 122 114 26 0.81 0.76 0.17
Bootstrap 150 123 90 36 0.82 0.60 0.24
Ohmyzsh 150 127 90 32 0.85 0.60 0.21
React 150 124 110 24 0.83 0.73 0.16
Vue 150 117 123 23 0.78 0.82 0.15
Overall 750 613 527 141 0.82 0.70 0.19
Table 16. Comparison with baselines based on the difficulty level of the questions
Level Questions Correct Accuracy
Our Approach MSRBot GPT-4o- Search-Preview Our Approach MSRBot GPT-4o- Search-Preview
1 200 174 123 101 0.87 0.62 0.51
2 395 304 311 30 0.77 0.79 0.08
3 155 135 93 10 0.87 0.60 0.06
Overall 750 613 527 141 0.82 0.70 0.19
@@
With few-shot chain-of-thought, the execution accuracy increases from 0.65 to 0.84 overall, and, for complex questions requiring two or more relationships, from 0.50 to 0.90. This implies that chain-of-thought can help answer complex questions requiring multiple relationships. The approach also outperformed the baselines, MSRBot (0.70) and GPT-4o-search-preview (0.19).
Figure 6. Distribution of correct responses across multiple runs for each project. The blue bar represents when all 5 answers are correct, the orange bar when 4 of the 5 answers are correct, the green bar when 3 of 5 are correct, the red bar when 2 of the 5 are correct, the purple bar when 1 of the 5 is correct and the brown bar when all 5 answers are incorrect.
5.5. RQ4: How do users perceive the usefulness of our approach in assisting them to answer repository-related questions?
Motivation. Building on the accuracy gains observed in RQ3, this research question examines whether these improvements translate into practical value for developers. Specifically, we examine whether users find our approach helpful in answering repository-related questions. To this end, we analyze task accuracy, completion time, and qualitative feedback, as these jointly capture both the efficiency gains and the subjective value users derive from the chatbot. This evaluation is critical for validating the approach from the end-user perspective, complementing quantitative performance metrics with insights into perceived utility.
Approach. To evaluate the usefulness of our approach with actual users, we conducted an online user study with 20 participants to assess the perceived usefulness of our approach. The study compared task performance when using our chatbot implementation of the approach against manual task execution. Participants first completed a set of repository-related tasks manually, then repeated similar tasks using the chatbot. We measured task accuracy and time taken, and collected qualitative feedback through a post-task questionnaire (see Section 4.5).
Table 17 shows the breakdown of the demographics of the study participants. Participants comprised 65.0% students and 35.0% industry practitioners, with the majority (55.0%) having over five years of Git experience. Most participants use Git daily (55.0%). Although the majority of our participants are students, this does not influence our findings for these reasons: most of the students (53.8% of student participants) have more than five years of experience using Git, 46.2% of the student population uses Git daily, and 53.2% of the student population uses Git weekly. Also, prior studies have shown that students with experience are a good proxy for industry experts, especially when the study is focused on emerging technology (Salman et al., 2015).
Results. Each of the 20 participants was tasked with completing 5 repository-related tasks manually and then performing them using the chatbot. The task completion rate and the correctness of the final answers for each participant are presented in Appendix F. The baseline (manual) approach led to the participants completing 69 out of the 100 tasks and providing a final answer, whereas, when using the chatbot, participants completed 91 tasks. Only three participants completed all five tasks manually, but 15 participants did so when using the chatbot, indicating the chatbot helped users finish more of their work.
Among the tasks that participants completed when working manually, they provided correct answers for 36% of the tasks, compared with 84% when using the chatbot. On average, participants solved approximately two tasks correctly without the assistance of our chatbot and more than four tasks correctly using the chatbot. No participant performed worse with the chatbot, and many improved from zero or one correct answer to four or five. These results indicate substantial gains in correctly obtaining information from the repository when using our approach.
Efficiency gains were also substantial. Across all five tasks, participants spent a median of 20.7 minutes when working manually, whereas using the chatbot required a median time of 10.3 minutes. Participants typically saved about 12 minutes per session, halving their effort. On a per-task basis, when using manual methods, participants took, on average, 4.5 minutes per answer, while using the chatbot reduced this to just over 2 minutes. Note that the 4.5-minute manual average reflects the 6-minute cap per task and is not an unconstrained effort time. Figure 7 shows the time participants spent on a task without using the chatbot and when using the chatbot.
The post-survey responses further underscore the chatbot’s perceived usefulness. Participants rated their agreement on a five-point scale and strongly agreed that the chatbot understood their questions (65% gave it the highest rating of 5), returned correct answers (50%), made the tasks easier (85%), and saved them significant time (80%). A majority also said they preferred using the chatbot over their usual methods (55% strongly agree) and would recommend it to others (60% strongly agree), with the remainder generally agreeing. The open-ended comments show a positive sentiment, with users calling the chatbot “amazing,” “very helpful and easy to use,” and noting that it solved problems they could not handle manually.
Overall, the results show that the chatbot not only boosts completion and accuracy but also substantially reduces effort. These benefits are clearly appreciated by the intended users.
Table 17. Demographics of Participants in the Survey
Category Experience Frequency %
Background Student 13 65
Industry Practitioner 7 35
Years of Experience with Git Tools 5+ years 11 55
3–5 years 4 20
1–2 years 5 25
Frequency of Using Git Tools Daily 11 55
Weekly 9 45
Total Participants 20
Figure 7. Boxplot of task completion times: participants consistently completed tasks faster with the chatbot than with manual methods.
RQ4 Summary:
Users perceived the chatbot as both accurate and efficient for answering their repository-related questions; it doubled the number of correct answers and saved time compared to other methods. Overall, participants found it helpful and preferable to other methods.
@@ 6. Discussions
In this study, we explored the synergy between large language models (LLMs) and knowledge graphs to enhance the accuracy of software engineering chatbots in answering software repository-related questions. Our findings demonstrate that synergizing LLMs and knowledge graphs can effectively improve the performance of LLM-based chatbots in answering software repositories-related questions. Beyond accuracy, our user study in RQ4 shows that these technical gains translate into practical value: participants completed more tasks, answered more correctly, did so in less time when using the chatbot, and reported high perceived usefulness.
We evaluated our approach on five popular open-source GitHub projects, with an accuracy range of 60%–75% in RQ1 and 80%–90% in RQ3. The difference in the accuracy of the approach on the projects is due to the non-deterministic nature of the LLM rather than the information in the projects. Our approach relies on the schema of the knowledge graph (see Figure 2), which is the same for all evaluated projects, to generate a query for retrieving the relevant data. Thus, if the LLM generated the same Cypher query for each question each time the question was executed, the accuracy would be constant across the repositories. This shows that the approach is generalizable across projects. Supplementing these findings, the baseline comparisons in RQ3 indicate that our method outperforms both MSRBot and the GPT-4o-search-preview model across projects and difficulty levels, reinforcing the benefit of combining knowledge and LLMs with CoT reasoning for repository Q&A.
We identified faulty reasoning as the prevalent challenge facing the approach, accounting for four of the limitations. This highlights the LLM’s challenges in accurately interpreting complex relationships and reasoning over the knowledge graph. In addressing the reasoning limitations, we introduced few-shot chain-of-thought prompting. This technique significantly improved the chatbot’s performance by 19% percentage points compared to the initial accuracy. In RQ4, this improvement aligned with observed usability gains: participants roughly halved their median task time with the chatbot while increasing correctness, suggesting that better reasoning support helps understand the user’s question and subsequently provide a correct answer.
@@
Our approach achieved results beyond the evaluation questions in Section 4. Our approach was able to answer ad-hoc queries that were not part of the original test set. For example, using lines of code as a metric for productivity (Maxwell et al., 1996; Cheng et al., 2022), we asked the question “Who added the most lines of code in December 2023” on the Vue project, it responded correctly by going through the right reasoning (See Appendix G). This highlights the promising potential of using LLMs with knowledge graphs to transform software engineering chatbots, making them more capable of handling a wide range of user queries. Together with the user study’s time savings and preference ratings, these observations suggest that the approach is not only accurate but also practically useful for everyday repository inquiries.
@@ 7.3. External Validity
External validity concerns the generalizability of our findings beyond this study. A first threat is generalizing beyond the evaluation questions. We evaluated on 150 questions covering 10 intents (Section 4). Although these questions cover different difficulty levels and intents, they may not encompass the full diversity of questions that users might pose in real-world scenarios. To mitigate this, we ran a task-based user study with 20 participants (Section 4.5) where users phrased tasks in their own words and worked both manually and with our chatbot. We also posed ad hoc questions outside the benchmark. These observations suggest some generalizability, but different question types could yield different results.
A second threat is the scope of our knowledge graph. The current schema models four entity types (Users, Commits, Issues, and Files), so questions that depend on non-modelled entities (e.g., pull requests, releases/tags, or CI/CD build events) cannot be answered. Although these four entities are critical to many repository-related questions, other entities could be added to the knowledge graph to answer more questions.
@@ 8. Conclusion
In this study, we investigated the synergy between large language models (LLMs) and knowledge graphs to improve the accuracy of software engineering chatbots in answering software repository-related questions. Our approach aimed to accurately answer natural language questions by generating Cypher queries to retrieve relevant repository data from the knowledge graph. Then use the retrieved information as context to generate a natural language response, making repository information accessible to both technical and non-technical stakeholders. We empirically evaluated our approach using five popular open-source GitHub repositories and a set of 20 questions curated from Abdellatif et al. (2020a) and categorized into three levels of difficulties. The findings demonstrated that LLMs, specifically the GPT-4o model, can answer repository-related questions by generating Cypher queries to retrieve accurate data from the knowledge graph. The initial accuracy of 65% achieved by our approach highlighted the potential limitation of synergizing LLMs with knowledge graphs. We manually investigated the instances where the approach failed to generate an accurate response and identified the faulty reasoning by the LLM as the predominant factor (80.5%) affecting the approach. We conducted further empirical evaluation if using few-shot chain-of-thought prompting can improve the accuracy. This technique significantly enhanced the reasoning ability of the LLM in our approach and improved the overall accuracy from 65% to 84%. There was a notable increase in the accuracy of the level 3 questions from 50% to 90%, signifying an improvement in the approach to handling complex queries. In addition to these gains, the baseline comparisons showed that combining LLMs with knowledge outperforms the intent-based approach, MSRBot, and GPT-4o-search-preview across projects and difficulty levels. Our findings highlight the integration of LLMs with knowledge graphs as a viable solution for making repository data accessible to both technical and non-technical stakeholders. The user study demonstrated that the technical improvements translate into practical benefits. In the study, the participants completed more tasks correctly and in less time with the chatbot than with their usual methods, and they reported high perceived usefulness. Also, our study highlights the importance of enhancing reasoning capabilities in LLMs. This opens avenues for further investigation in this direction. In this study, we focus on Git metadata (Users, Commits, Issues, Files) and do not answer questions that require program analysis. Future work can extend the repository knowledge graph with code entities and program-analysis edges and align the code-level entities with the existing metadata to answer questions about code comprehension.